Sean is buying salad and pizza for a company lunch. Suppose that a bowl of salad costs $5.00, and a slice of pizza costs $3.00.

Let E

E be the amount in dollars that Sean spends on salad and pizza. If Sean buys S

S bowls of salad and P

P slices of pizza, then the total amount of money he spends (E

E) can be represented by the equation    .


Now rearrange the equation you wrote above so that P

P is written in terms of E

E and S

S. The quantity of pizza he buys can be represented by the equation    .


Suppose Sean has $30.00 to spend on salad and pizza; that is, E=$30.00

E=$30.00.


1
Expert's answer
2021-09-06T18:53:31-0400

Assuming individual-individual C exists. Sean is purchasing Salad (PS) costing 5 dollars as well as pizza (PP) costing 3 dollars. In this case, E is let to be cumulative amount spent on P and S by C.




Absolute spending is similar to amount of good P and S expenses.

"E = 5.00 S + 3P"


In S and E, P equation is composed.


"3P = E - 5.00S"


"P = \frac{E-5.00S}{3}"

The needed equation is therefore;


"P = \frac{E-5.00S}{3}"

Individual C will therefore have $30 spending on goods E.